# Strong and noise-tolerant entanglement in dissipative optomechanics

### v3

@@ IV Covariance matrix
With steady-state values in hand, we further expand each operator $\mathcal{O}$ around its steady state, i.e., $\mathcal{O}=\delta\mathcal{O}+\mathcal{O}_{s}$. Substituting this expression back into Eq. (III), we then obtain a set of equations for fluctuations ($\delta\mathcal{O}$). For the strong driving field ($n_{s}\gg 1$), the higher-order fluctuation terms can be safely neglected Vitali et al. (2007), corresponding to linearization. As a result, the dynamics for fluctuations can be written as
@@
$\dot{\delta a}=$ $-(\kappa_{s}+i\Delta_{s})\delta a+(\frac{G_{\kappa}+iG_{\omega}}{\sqrt{2}}-\frac{g_{\kappa}\mathcal{E}}{\sqrt{2\kappa_{a}}})\delta x+x_{s}^{\prime}a_{\rm in},$
@@
where $u^{T}(t)=(\delta x,\delta p,\delta x_{a},\delta p_{a})$ is the vector of fluctuation operators, $n(t)$ is the vector of noise operator, given by $n^{T}(t)=(0,\xi+\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}p_{\rm in},x_{s}^{\prime}x_{\rm in},x_{s}^{\prime}p_{\rm in})$, and the drift matrix $A$ is
$A=\left(\begin{array}[]{cccc}0&\omega_{m}&0&0\\ -\omega_{m}&-\gamma_{m}&G_{\omega}+\frac{G_{\kappa}\Delta_{s}}{\sqrt{2\kappa_{a}}x_{s}^{\prime}}&-\frac{G_{\kappa}\kappa_{s}}{\sqrt{2\kappa_{a}}x_{s}^{\prime}}\\ G_{\kappa}-\frac{2G_{\kappa}\kappa_{s}}{\sqrt{2\kappa_{a}}x_{s}^{\prime}}&0&-\kappa_{s}&\Delta_{s}\\ G_{\omega}-\frac{2G_{\kappa}\Delta_{s}}{\sqrt{2\kappa_{a}}x_{s}^{\prime}}&0&-\Delta_{s}&-\kappa_{s}\\ \end{array}\right).$ (16)
@@
Since the phonon noise $\xi$ and the photon noise $a_{\rm in}$ are zero-mean quantum Gaussian noises and the dynamics is linearized [see Eq. (IV)], the quantum steady state for fluctuations is a zero-mean bipartite Gaussian state, fully described by a $4\times 4$ correlation matrix ${V}$, where $V_{ij}=\langle u_{i}(t)u_{j}(t^{\prime})+u_{j}(t^{\prime})u_{i}(t)\rangle/2$ with $i,j=1,2,3,4$. The $V$ is obtained by solving the Lyapunov equation $AV+VA^{T}=-D$, where $D$ is the diffusion matrix, defined by $\langle n_{i}(t)n_{j}(t^{\prime})+n_{j}(t^{\prime})n_{i}(t)\rangle/2=D_{ij}\delta(t-t^{\prime})$. More specifically, $D$ can be expressed as
$D=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&\gamma_{m}(2n_{\rm th}+1)+\frac{G_{\kappa}^{2}}{4\kappa_{a}}&0&\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}\frac{x_{s}^{\prime}}{2}\\ 0&0&\frac{x_{s}^{\prime,2}}{2}&0\\ 0&\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}\frac{x_{s}^{\prime}}{2}&0&\frac{x_{s}^{\prime,2}}{2}\\ \end{array}\right).$ (17)
Compared to coherent coupling case, dissipative coupling strength appears in the off-diagonal elements of the diffusion matrix, playing an crucial role in generation of bipartite entanglement below.
@@ V.1 Coherent coupling ($g_{\kappa}=0$)
In the coherent coupling case ($g_{\kappa}=0$), we present the entanglement measure $E_{N}$ as a function of both the coherent coupling strength $g_{\omega}$ and the normalized effective cavity detuning $\Delta_{s}/\omega_{m}$ in Fig. 3(a) to explore the conditions for optimal optomechanical entanglement at a bath temperature of $T=0.4$ K. The results reveal that the generated entanglement remains relatively weak compared to that typically observed in the conventional optomechanical system Vitali et al. (2007). This is because the stronger coupling strength will lead to system unstable. To numerically ensure system stable, the coupling strength is required to be weak, giving rise to weak entanglement. Nonetheless, optimal entanglement can still be achieved around the point ($\Delta_{s}/\omega_{m},g_{\omega}/2\pi)\approx(2,3.1)$, as marked by the green star, indicating that even under these constraints, entanglement can be optimized.
@@
To further investigate the coherent coupling effect on entanglement, we plot $E_{N}$ as a function of $g_{\omega}$ in Fig. 3(b), where $\Delta_{s}\approx 2\omega_{m}$ is fixed. It is obviously shown that $E_{N}$ increases with $g_{\omega}$. Once up to its maximum, $E_{N}$ will have a sharp reduction by further increasing $g_{\omega}$. Such abrupt behavior corresponds to the transition from the stability to instability, as shown by the pink-shaded region. Additionally, when the bath temperature increases to $T=4$ K [see the blue dashed curve], the coupling for generation of quantum entanglement is required to be larger and $E_{N}$ has a significant reduction. This indicates that quantum entanglement can be affected by the bath temperature. To clearly evaluate the effect of bath temperature, we plot $E_{N}$ as a function of the bath temperature in Fig. 3(c), where $g_{\omega}/2\pi=3.1$ Hz is fixed to obtain optimal entanglement. As the temperature increases, $E_{N}$ gradually declines. When the bath temperature increases to $T\approx 8.5$ K, $E_{N}=0$, that is, quantum entanglment vanishes and the system becomes separable.
@@ V.2 Dissipative coupling ($g_{\omega}=0$)
In dissipative coupling case ($g_{\omega}=0$), we observe that the optimal entanglement can be significantly stronger than in the coherent coupling case, even surpassing the entanglement achieved in conventional optomechanical system (i.e., $E_{N}=0.35$Vitali et al. (2007). This is demonstrated in Fig. 4(a), where $E_{N}$ is plotted as a function of both the dissipative coupling strength ($g_{\kappa}$) and the normalized effective cavity detuning $\Delta_{s}/\omega_{m}$. The optimal entanglement is found at approximately ($\Delta_{s}/\omega_{m},g_{\kappa}/2\pi)\approx(0.1,20)$, as indicated by the green star. In contrast to the coherent coupling case [see Fig. 2(a)], where the optimal entanglement occurs near the mechanical red sideband at $\Delta_{s}/\omega_{m}=2$, the optimal entanglement in the dissipative coupling case is obtained when the cavity is nearly resonant with the driving field, specifically at $\Delta_{s}/\omega_{m}=0.1$, which is far removed from the mechanical red sideband.
@@
Furthermore, in Fig. 4(b), we investigate the behavior of the optomechanical entanglement as a function of the dissipative coupling strength, while fixing $\Delta_{s}=0.1\omega_{m}$. The resulting behavior of $E_{N}$ mirrors that observed in the coherent coupling case (see the red curve). The abrupt drop in entanglement marks the transition of the system from a stable to an unstable region, as indicated by the pink area. Notably, a relatively large dissipative coupling strength is required to achieve the optimal entanglement. This suggests that in the regime dominated by dissipative coupling, stronger interactions are necessary to generate significant quantum effects. As the bath temperature increases, for instance to $T=4$ K, the entanglement is reduced (see the blue curve). To explore the influence of bath temperature on optomechanical entanglement, we fix $g_{\kappa}/2\pi=19$ Hz and present the results in Fig. 4(c). We observe a sharp decrease in entanglement as the system temperature increases from 0 K to 5 K. However, beyond this point, the entanglement becomes more tolerant to further temperature increases, suggesting that the dissipative coupling becomes more robust to thermal noise at higher temperatures. Remarkably, entanglement persists even up to 25 K, This indicates that strong, noise-resistant entanglement can be generated through dissipative optomechanical coupling, opening new possibilities for quantum information processing and communication in noisy
@@ V.3 Cooperative coupling ($g_{\omega}\neq 0$ and $g_{\kappa}\neq 0$)
To examine the effect of the cooperative coupling (i.e., coherent and dissipative couplings) on optomechanical entanglement, we set the coherent coupling strength to be $g_{\omega}/2\pi=3$ Hz, corresponding to the optimal entanglement condition at $\Delta_{s}/\omega_{m}=2$ in the dispersive coupling case [refer to Fig. 3(a)]. By progressively increasing the dissipative coupling strength, we observe a gradual decline in the entanglement, as shown in Fig. 5(a). The reduction in entanglement possibly arises from the interference between the coherent and dissipative couplings. When the bath temperature increases to $T=4$ K, it is observed that the optomechancial entanglement is further suppressed, as depicted by the blue dashed curve in Fig. 5(b). Moreover, the larger $g_{\kappa}$ is not benificial to entanglement generation. This can also be revealed by Fig. 5(c), where $g_{\kappa}/g_{\omega}=0.01$ (the red curve) and $g_{\kappa}/g_{\omega}=0.1$ (the blue curve) are taken . For $g_{\kappa}/g_{\omega}=0.1$, the survival temperature of the entanglement is about $5$ K. But for $g_{\kappa}/g_{\omega}=0.01$, the survival temperature can be up to $8$ K.
@@ VI Conclusion
In a summary, we investigate quantum entanglement in a dissipative optomechanical system, realized by a fixed mirror and a Michelson–Sagnac interferometer with a membrane. This system allows the dissipative and coherent couplings to be independently tuned on and off by adjusting the membrane position or the reflectivity (transmissivity) of the beam splitter, enabling us to explore quantum entanglement in three distinct scenarios: coherent coupling case, dissipative case, and cooperative case. Our results reveals that, in the coherent coupling case, the steady-state mechanical displacement exhibits a nonlinear dependence on the driving power, whereas in the dissipative coupling case, the displacement shows a linear dependence. Furthermore, entanglement generated through dissipative coupling is not only significantly stronger but also more noise-tolerant compared to that achieved via coherent coupling. When both coupling mechanisms coexist, quantum entanglement is reduced, likely due to the interference between the two coupling interactions. These findings suggest that dissipative coupling may be a promising mechanism to engineer strong and noise-tolerant quantum entanglement.

### v4

@@ IV Covariance matrix
Apart from the steady-state values, a set of equations related to the fluctuations $\delta\mathcal{O}$ can be obtained when the expression of $\mathcal{O}=\mathcal{O}_{s}+\delta\mathcal{O}$ is substituted back into Eq. (III). Under the strong driving field ($n_{s}\gg 1$), the higher-order fluctuation terms can be safely neglected Vitali et al. (2007), which corresponds to the linearization of the system dynamics. As a result, the dynamics for the fluctuations can be written as
@@
$\dot{\delta a}=$ $-(\kappa_{s}+i\Delta_{s})\delta a+(\frac{G_{\kappa}+iG_{\omega}}{\sqrt{2}}-\frac{g_{\kappa}\mathcal{E}}{\sqrt{2\kappa_{a}}})\delta x+{\Gamma_{s}}a_{\rm in},$
@@
where $u^{T}(t)=(\delta x,\delta p,\delta x_{a},\delta p_{a})$ is the vector of fluctuation operators, $n(t)$ is the vector of noise operator, given by $n^{T}(t)=(0,\xi+\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}p_{\rm in},{\Gamma_{s}}x_{\rm in},{\Gamma_{s}}p_{\rm in})$, and the drift matrix $A$ is
$A=\left(\begin{array}[]{cccc}0&\omega_{m}&0&0\\ -\omega_{m}&-\gamma_{m}&G_{\omega}+\frac{G_{\kappa}\Delta_{s}}{\sqrt{2\kappa_{a}}{\Gamma_{s}}}&-\frac{G_{\kappa}\kappa_{s}}{\sqrt{2\kappa_{a}}{\Gamma_{s}}}\\ G_{\kappa}-\frac{2G_{\kappa}\kappa_{s}}{\sqrt{2\kappa_{a}}{\Gamma_{s}}}&0&-\kappa_{s}&\Delta_{s}\\ G_{\omega}-\frac{2G_{\kappa}\Delta_{s}}{\sqrt{2\kappa_{a}}{\Gamma_{s}}}&0&-\Delta_{s}&-\kappa_{s}\\ \end{array}\right).$ (17)
@@
Since the phonon noise $\xi$ and the photon noise $a_{\rm in}$ are zero-mean quantum Gaussian noises and the dynamics is linearized [see Eq. (IV)], the quantum steady state for fluctuations is a zero-mean bipartite Gaussian state, fully described by a $4\times 4$ correlation matrix ${V}$, where $V_{ij}=\langle u_{i}(\infty)u_{j}(\infty)+u_{j}(\infty)u_{i}(\infty)\rangle/2$ with $i,j=1,2,3,4$. The $V$ is obtained by solving the Lyapunov equation $AV+VA^{T}=-D$, where $D$ is the diffusion matrix, defined by $\langle n_{i}(t)n_{j}(t^{\prime})+n_{j}(t^{\prime})n_{i}(t)\rangle/2=D_{ij}\delta(t-t^{\prime})$. More specifically, $D$ can be expressed as
$D=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&\gamma_{m}(2n_{\rm th}+1)+\frac{G_{\kappa}^{2}}{4\kappa_{a}}&0&\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}\frac{{\Gamma_{s}}}{2}\\ 0&0&\frac{{\Gamma_{s}^{2}}}{2}&0\\ 0&\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}\frac{{\Gamma_{s}}}{2}&0&\frac{{\Gamma_{s}^{2}}}{2}\\ \end{array}\right).$ (18)
Compared to the coherent coupling case, dissipative coupling strength appears in the off-diagonal elements of the diffusion matrix, playing a crucial role in the generation of bipartite entanglement below.
@@ V.1 Coherent coupling ($g_{\kappa}=0$)
In the coherent coupling case ($g_{\kappa}=0$), we present the entanglement measure $E_{N}$ as a function of both the coherent coupling strength $g_{\omega}$ and the normalized effective cavity detuning $\Delta_{s}/\omega_{m}$ in Fig. 3(a) to explore the conditions for optimal optomechanical entanglement at a bath temperature of $T=0.4$ K. The results reveal that the generated entanglement remains relatively weak compared to that typically observed in the conventional optomechanical system Vitali et al. (2007). This is because a weaker coupling strength is necessary to maintain system stability, which in turn leads to weaker entanglement, while a stronger coupling strength can induce instability. Nonetheless, optimal entanglement can still be achieved around the point ($\Delta_{s}/\omega_{m},g_{\omega}/2\pi)\approx(2,3.1)$, as marked by the green star, indicating that even under these constraints, entanglement can be optimized.
@@
To further investigate the coherent coupling effect on entanglement, we plot $E_{N}$ as a function of $g_{\omega}$ in Fig. 3(b), where $\Delta_{s}\approx 2\omega_{m}$ is fixed. It is obviously shown that $E_{N}$ increases with $g_{\omega}$. Once up to its maximum, $E_{N}$ will have a sharp reduction by further increasing $g_{\omega}$. Such abrupt behavior corresponds to the transition from stability to instability, as shown by the pink-shaded region. Additionally, when the bath temperature increases to $T=4$ K [see the blue dashed curve], the coupling for generation of quantum entanglement is required to be larger and $E_{N}$ has a significant reduction. This indicates that quantum entanglement can be affected by the bath temperature. To clearly evaluate the effect of bath temperature, we plot $E_{N}$ as a function of the bath temperature in Fig. 3(c), where $g_{\omega}/2\pi=3.1$ Hz is fixed to obtain optimal entanglement. As the temperature increases, $E_{N}$ gradually declines. When the bath temperature increases to $T\approx 8.5$ K, $E_{N}=0$, which indicates that quantum entanglement vanishes and the system becomes separable.
@@ V.2 Dissipative coupling ($g_{\omega}=0$)
In dissipative coupling case ($g_{\omega}=0$), we observe that the optimal entanglement can be significantly stronger than in the coherent coupling case, even surpassing the entanglement achieved in conventional optomechanical systems (i.e., $E_{N}=0.35$Vitali et al. (2007). This is demonstrated in Fig. 4(a), where $E_{N}$ is plotted as a function of both the dissipative coupling strength ($g_{\kappa}$) and the normalized effective cavity detuning $\Delta_{s}/\omega_{m}$. The optimal entanglement is found at approximately ($\Delta_{s}/\omega_{m},g_{\kappa}/2\pi)\approx(0.1,20)$, as indicated by the green star. In contrast to the coherent coupling case [see Fig. 2(a)], where the optimal entanglement occurs near the mechanical red sideband at $\Delta_{s}/\omega_{m}=2$, the optimal entanglement in the dissipative coupling case is obtained when the cavity is nearly resonant with the driving field, specifically at $\Delta_{s}/\omega_{m}=0.1$, which is far removed from the mechanical red sideband.
@@
Furthermore, in Fig. 4(b), we investigate the behavior of the optomechanical entanglement as a function of the dissipative coupling strength, while fixing $\Delta_{s}=0.1\omega_{m}$. The resulting behavior of $E_{N}$ mirrors that observed in the coherent coupling case (see the red curve). The abrupt drop in entanglement marks the transition of the system from a stable to an unstable region, as indicated by the pink area. Notably, a relatively large dissipative coupling strength is required to achieve the optimal entanglement. This suggests that in the regime dominated by dissipative coupling, stronger interactions are necessary to generate significant quantum effects. As the bath temperature increases, for instance to $T=4$ K, the entanglement is reduced (see the blue curve). To explore the influence of bath temperature on optomechanical entanglement, we fix $g_{\kappa}/2\pi=19$ Hz and present the results in Fig. 4(c). We observe a sharp decrease in entanglement as the system temperature increases from 00 K to $5$ K. However, beyond this point, the entanglement becomes more tolerant to further temperature increases, suggesting that the dissipative coupling becomes more robust to thermal noise at higher temperatures. Remarkably, entanglement persists even up to $25$ K. This indicates that strong, noise-resistant entanglement can be generated through dissipative optomechanical coupling, opening new ways for quantum information processing and communication in dissipative systems.
@@ V.3 Cooperative coupling ($g_{\omega}\neq 0$ and $g_{\kappa}\neq 0$)
To examine the effect of the cooperative coupling (i.e., coherent and dissipative couplings) on optomechanical entanglement, we set the coherent coupling strength to be $g_{\omega}/2\pi=3$ Hz, corresponding to the optimal entanglement condition at $\Delta_{s}/\omega_{m}=2$ in the dispersive coupling case [refer to Fig. 3(a)]. By progressively increasing the dissipative coupling strength, we observe a gradual decline in the entanglement, as shown in Fig. 5(a). This degradation of entanglement results from the destructive interference between the coherent and dissipative coupling mechanisms. When the bath temperature increases to $T=4$ K, the entanglement is further reduced, as depicted by the blue dashed curve in Fig. 5(b). Furthermore, a larger $g_{\kappa}$ is found to be detrimental to the generation of entanglement. This is further illustrated in Fig. 5(c), where $g_{\kappa}/g_{\omega}=0.01$ (the red curve) and $g_{\kappa}/g_{\omega}=0.1$ (the blue curve) are considered. For $g_{\kappa}/g_{\omega}=0.1$, the survival temperature of the entanglement is about $5$ K, whereas for $g_{\kappa}/g_{\omega}=0.01$, the survival temperature can reach up to $8$ K.
@@ VI Conclusion
Before concluding, we find it worthwhile to briefly discuss the conditions under which the coherent and dissipative coupling strengths, $g_{\omega}$ and $g_{\kappa}$, can become comparable or even equal, particularly in the context of implementing either purely coherent or purely dissipative coupling mechanisms. In our setup, based on the Michelson-Sagnac interferometer configuration Xuereb et al. (2011); Tarabrin et al. (2013); Sawadsky et al. (2015); Friedrich et al. (2011), the single-photon coherent and dissipative coupling strengths can be independently tuned by adjusting the reflectivity of the beam splitter and the position of the membrane. This tunability allows access to both purely dispersive coupling ($g_{\kappa}=0$) and purely dissipative coupling ($g_{\omega}=0$). However, we emphasize that achieving equal values of $g_{\omega}$ and $g_{\kappa}$ under their respective coupling regimes involves different levels of experimental complexity. For instance, realizing a purely dissipative interaction typically demands more precise interferometric alignment and control compared to the conventional dispersive case. As a result, achieving equal values of $g_{\omega}$ in a purely coherent optomechanical system and $g_{\kappa}$ in a purely dissipative optomechanical system is generally challenging. Fortunately, under strong driving conditions, the optomechanical system can be linearized, resulting in effective coupling strengths that are proportional to the amplitude of the driving field [see Eq. (IV)]. This enables the use of two independent MSI setups, one exhibiting purely dispersive coupling and the other purely dissipative coupling, each driven by an appropriately chosen strong optical field. Although the intrinsic coupling strengths $g_{\omega}$ and $g_{\kappa}$ may differ, by tailoring the driving field amplitudes, one can achieve comparable or even equal effective coupling strengths, $G_{\omega}$ and $G_{\kappa}$. This approach ensures a fair comparison in the investigation of quantum entanglement, as presented in Figs. 3 and 4, despite the inherent differences in the physical realization of the coupling mechanisms.
In summary, we investigate quantum entanglement in a dissipative optomechanical system, realized by a fixed mirror and a Michelson-Sagnac interferometer with a membrane. This system allows the dissipative and coherent couplings to be independently tuned on and off by adjusting the membrane position or the reflectivity (transmissivity) of the beam splitter, enabling us to explore quantum entanglement in three distinct scenarios: coherent coupling, dissipative coupling, and cooperative coupling. Our results reveal that, in the coherent coupling case, the steady-state mechanical displacement exhibits a nonlinear dependence on the driving power, while in the dissipative coupling case, the displacement shows a linear response. Furthermore, entanglement generated through dissipative coupling is not only significantly stronger but also more noise-tolerant compared to that achieved via coherent coupling. However, when both coupling mechanisms coexist, quantum entanglement is suppressed due to the destructive interference between the coherent and dissipative interactions. These findings suggest that dissipative coupling offers a promising mechanism to engineer strong and noise-tolerant quantum entanglement.